OpenStudy (anonymous):

If the volume of the sphere is 18 cm cubed, what is the radius?

OpenStudy (anonymous):

oh lord

OpenStudy (anonymous):

hey hottie

OpenStudy (anonymous):

Do you have the equation for the volume of a sphere?

OpenStudy (anonymous):

yes it is; 4/3*pi*radius cubed

OpenStudy (anonymous):

So just solve for R by moving 3/(4*pi) to the other side of the equation so you get R^3 = (3*18)/(4*pi) so R=1.63